### v1

# Robust Bayesian Scene Reconstruction by Leveraging Retrieval-Augmented Priors

Abstract
Constructing 3D representations of object geometry is critical for many downstream manipulation tasks. These representations must be built from potentially noisy partial observations. In this work we focus on the problem of reconstructing a multi-object scene from a single RGBD image. Current deep learning approaches to this problem can be brittle to noisy real world observations and out-of-distribution objects. Other approaches that do not rely on training data cannot accurately infer the backside of objects. We propose BRRP, a reconstruction method that can leverage preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction. In order to make our method more efficient, we introduce the concept of retrieval-augmented prior, where we retrieve relevant components of our prior distribution during inference. Our method produces a distribution over object shape that can be used for reconstruction or measuring uncertainty. We evaluate our method in both procedurally generated scenes and in real world scenes. We show our method is more robust than a deep learning approach while being more accurate than a method with an uninformative prior.
I Introduction
The ability to construct internal representations of its operating environment is key for robot autonomy. These representations need to be particularly fine-grained for robotic manipulation, which often requires closely interacting with and avoiding objects. These interactions make it necessary for robots to develop an understanding of the geometry within their vicinity. Explicit 3D representations of the geometry of the scene are often required for the robust usage of downstream grasping and motion planning algorithms. These representations must be built from observations that are both noisy and only contain partial information of the scene. In our case, we focus on the problem of building a 3D representation of multi-object scenes from a single RGBD camera image.
One approach to this problem is to build some type of neural network that predicts the full geometry of an object given a partial view. Such approaches are able to use existing mesh datasets to more accurately infer the backside of objects. Unfortunately, these approaches also tend to have a number of problems when used on the real world depth cameras. The presence of unknown objects, significant occlusion, noisy point clouds, or inaccurate segmentation prevent these methods from being reliably deployed in less-structured environments.
Most deep learning methods for 3D reconstruction also often lack a sense of uncertainty about the shape of the object. Such uncertainty has been shown to be useful in the downstream task of robotic grasping [1, 2]. Such uncertainty can also be useful in developing safe motion plans or during active learning.
Another common approach to building full 3D representations from a partial view is to simply build such a representation in a way that does not rely on prior information from mesh datasets. A common example of this is the Gaussian Process Implicit Surface [3] model. A more recent example is V-PRISM [4], which probabilistically maps tabletop scenes without using prior information. These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution. However, this also means that they don’t always accurately reconstruct the backside of objects.
Figure 1: Our method, BRRP, (a), (b) takes an input segmented RGBD image, and (c) retrieves objects to use as a prior, which allow it to (d) reconstruct the scene as well as (e) capture principled uncertainty about object shape.
In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors. Our method is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects. In order to improve efficiency, we introduce the idea of a retrieval-augmented prior, inspired by [5], where we retrieve relevant components of our prior distribution based on classification results. We refer to our method as BRRP (pronounced ’burp’), which stands for Bayesian Reconstruction with Retrieval-augmented Priors. BRRP starts with an observed RGBD image with corresponding instance segmentations. Then, we compute an identification result that predicts which objects from our database should be retrieved to use as a prior during reconstruction. We then negatively sample the observation and use Stein Variational Gradient Descent (SVGD) [6] to infer a posterior distribution over object shapes. This distribution is defined as a set of Hilbert Maps [7] for each object. Because we solve for a distribution over weights, we can recover principled uncertainty about the shape of each object. In practice we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant. An example scene and reconstruction can be seen in Figure 1.
We conduct experiments on BRRP in both procedurally generated scenes and in the real world. We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects on the generated scenes. We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera. We also qualitatively examine the uncertainty from our method’s reconstructions.
Our three primary contributions can be summarized as follows:
  1. 1.
    The formulation of retrieval-augmented priors for Bayesian inference.
  2. 2.
    A formulation for a prior over Hilbert maps with pose and scale invariance
  3. 3.
    A novel, robust Bayesian scene reconstruction method that utilizes prior information from existing mesh datasets
Our paper has the following organization. We overview related works in Section II. In Section III we cover the necessary mathematical preliminaries for our method. In Section IV, we introduce the concept of retrieval-augmented priors. Our method, BRRP, is introduced in Section V. Then, we perform experiments in Section VI before a conclusion in Section VII
II Related Works
Explicit 3D Representations. There are many different ways of representing 3D geometry of a scene. In the mapping literature, techniques such as truncated signed distance functions [8] build voxelized representations of an environment. Hilbert maps [7], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space. Hilbert maps representations have also been extended to Bayesian Hilbert maps of various forms [9, 10, 11]. Neural implicit functions have also been used to represent continuous 3D geometry [12, 13, 14]. Other representations are built using differentiable rendering and combining multiple views. Neural radiance fields [15] are an example of this, in which a neural network maps 3D position to density and color. 3D Gaussian splatting [16] does a similar thing but with a set of Gaussians instead of a neural network. Another representation that has been studied is super quadrics [17].
3D Reconstruction with Deep Learning. Many methods have been proposed as ways to leverage deep learning to reconstruct scenes or objects. While some methods aim to predict object shape from RGB data only [18, 19, 20, 21], we instead focus on using depth measurements during reconstruction. DeepSDF [12] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function. In the context of robotics, [22] extends DeepSDF to have uncertainty-awareness. Other work, such as occupancy networks [13] or PointSDF [23] try to directly predict such a latent code without inference-time optimization. Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [24, 25]. Language is also used during reconstruction in [26] and [27]. Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [28]. While these work typically focus on single object scenes, other work focuses on reconstruction scenes with multiple, occluding objects. For example, [29] learns a reconstruction from a voxel representation with different channels to account for occlusion. Another method uses silhouettes to refine the initial reconstructions [30]. In practice these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras. Inaccurate segmentation can also be a problem for many of these methods as well.
3D Reconstruction without Deep Learning. There are also many approaches to perform probabilistic 3D reconstruction without deep learning. Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [31]. Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [3]. While there is an extension of GPIS to a slightly more informative prior [32], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors. In our work, we derive our prior from pre-existing mesh datasets. V-PRISM [4] is another method that probabilistically maps the scene using a multiclass framing.
Using Reconstructions in Manipulation. 3D reconstruction methods have seen extensive use in manipulation. In [23], PointSDF provides collision constraints during grasping. PointSDF is also utilized in [33], where tactile sensors are used along with the reconstruction during grasping. A learning-based voxel representation is used for grasping in [34]. Neural shape completion is also used during the anthropomorphic grasping pipeline proposed in [35]. GPIS is also a common representation for manipulation applications. Some recent work has utilized the uncertainty from GPIS representations during grasp selection [1, 2]. We believe BRRP provides principled uncertainty measurements that can similarly be utilized in downstream manipulation tasks.
III Background
III-A Hilbert Maps
A Hilbert map [7] is a continuous occupancy map. It represents the environment by a continuous function that is defined by a linear function of a fixed feature transform. Typically this feature transform is induced by a set of hinge points, $\{\mathbf{h}_{1},...,\mathbf{h}_{H}\}\subset\mathbb{R}^{3}$ and a translation-invariant kernel $k(d)$. The transform is then defined as:
$\phi(\mathbf{x})=[1,k(\mathbf{x}-\mathbf{h}_{1}),...,k(\mathbf{x}-\mathbf{h}_{H})]^{\top}.$
Typically, a Gaussian kernel is used and hinge points are placed in an evenly-spaced grid. An occupancy map can then be defined by a single weight vector, $\mathbf{w}\in\mathbb{R}^{H+1}$, as such:
$m(\mathbf{x})=\sigma(\mathbf{w}^{\top}\phi(\mathbf{x})).$
Figure 2: (a) A hinge point feature transform induced by a set of hinge points is used by Hilbert maps [7]; (b) these maps are built by first sampling negative samples along the unoccupied portions of the camera ray.
In order to recover the weight vector corresponding to a given depth observation, negative sampling is performed along the unoccupied portions of the depth rays. These negative samples are assigned a label of unoccupied and the points at the end of the ray are labeled as occupied. Then, stochastic gradient descent (SGD) is performed on the binary cross entropy of the negative samples and terminal points of the ray. The binary cross entropy measures the likelihood of the samples, and is defined as:
$\text{BCE}(y,\mathbf{w}^{\top}\phi(\mathbf{x}))=\begin{cases}-\ln\left[\sigma\left(\mathbf{w}^{\top}\phi(\mathbf{x})\right)\right],&y=1\\ -\ln\left[1-\sigma\left(\mathbf{w}^{\top}\phi(\mathbf{x})\right)\right],&y=0\end{cases}$ (1)
Figure 2 shows an illustration of both the hinge point feature transform and the negative samples used during Hilbert map construction.
Hilbert maps have previously been extended to Bayesian Hilbert maps, where a distribution over weights is modeled as a multivariate Gaussian [9, 10, 11]. There is also a multiclass variant that defines a weight matrix with Gaussian rows described in [4]. In this work, we adopt the Hilbert map representation, but model the distribution over weights as a collection of particles. This allows our method the capability to capture irregular, non-Gaussian posterior distributions.
III-B Stein Variational Gradient Descent
Figure 3: Overview of BRRP method. We begin with a segmented RGBD image and (a) feed cropped images of each segment into CLIP to get object probabilities (Section V-B) Then, we retrieve and (b) register the the top-k objects in the prior. This gives us a set of registered prior samples (Section V-A). We also (c) compute negative samples based on the observed segmented point cloud (Section V-C). Finally, (d) we run SVGD optimization to recover a posterior distribution over Hilbert map weights (Section V-D). We can use this distribution to both reconstruct the scene as well as measure uncertainty.
Stein Variational Gradient Descent (SVGD) [6] is a algorithm for variational inference that closely resembles gradient descent. The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$. Usually, this takes the form of an optimization problem over the Kullback-Leibler (KL) divergence:
$q^{*}=\text{arg}\min_{q\in\mathcal{Q}}\mathbb{KL}(q\|p).$
SVGD aims to iteratively transform $q$ in descent directions of the KL divergence in a $d$-dimensional reproducing kernel Hilbert space (RKHS), $\mathcal{H}^{d}$. Because this Hilbert space is a space of functions, a descent direction requires deriving the functional gradient of our KL divergence objective.
Theorem 1
From [6]. Let $T(\mathbf{x})=\mathbf{x}+f(\mathbf{x})$, where $f\in\mathcal{H}^{d}$ and $q_{[T]}$ is the density of random variable $\mathbf{z}=T(\mathbf{x})$ when $\mathbf{x}\sim q$. Then
$\nabla_{f}\mathbb{KL}(q_{[T]}\|p)|_{f=0}=-g^{*}_{q,p},$
where $g^{*}_{q,p}=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[k(\mathbf{x},\cdot)\nabla_{\mathbf{x}}\ln p(\mathbf{x})+\nabla_{\mathbf{x}}k(\mathbf{x},\cdot)]$.
In SVGD, $q$ is approximated by a set of particles $\mathbf{x}^{(0)}_{1},...,\mathbf{x}^{(0)}_{P}\sim q(\mathbf{x})$. This can be used to approximate the gradient in Theorem 1 with $\hat{g}^{*}$:
$\hat{g}^{*}(\mathbf{x})=\frac{1}{P}\sum_{i=1}^{P}k(\mathbf{x}_{i},\mathbf{x})\nabla_{\mathbf{x_{i}}}\ln p(\mathbf{x}_{i})+\nabla_{\mathbf{x_{i}}}k(\mathbf{x},\cdot).$ (2)
The particles can then be iteratively updated according to Equation 2:
$\mathbf{x}^{(t+1)}_{i}=\mathbf{x}^{(t)}_{i}+\epsilon\hat{g}^{*}(\mathbf{x}^{(t)}_{i})$
The result of these iterations is that the set of particles converges to an approximation of the target distribution $p$. Importantly, Equation 2 only relies on the gradient of the log of $p$, which means we can perform variational inference to an unnormalized distribution. Such unnormalized distributions are commonplace in many Bayesian inference problems, including ours.
IV Retrieval-Augmented Priors
Retrieval-augmented generation [5] was originally introduced in the context of improving language generation. The work has served as inspiration for an approach to affordance-prediction in [36]. In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior distributions during Bayesian inference.
To motivate retrieval-augmented priors, consider the problem of Bayesian inference with a mixture model acting as the prior distribution. Given some data, we would like to infer a posterior distribution over hypotheses. If we have a mixture model as a prior distribution, then:
$P(H|D)\propto P(D|H)\sum_{c=1}^{C}P(H|c).$ (3)
If our prior distribution has a lot of components, it may be inefficient to fully evaluate. This could be a serious problem for algorithms like SVGD, which requires iteratively computing the gradient of both the likelihood and prior. Inspired by [5], the insight behind retrieval-augmented priors is to determine which subset of the prior distribution components to retrieve and use given some detection result $R$. Conditioning on this detection result, we have a new posterior distribution, $P(H|D,R)$. Making an independence assumption,
$P(H|D,R)\propto P(D|H)\cdot\mathbb{E}_{c\sim P(c|R)}[P(H|c)].$
Comparing to Equation 3, the expectation now replaces the true prior. Then, we can use a top-$k$ approximation for the expectation:
$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$
This means that we only need to evaluate a subset of the prior distribution components.
V The BRRP Method
Our method takes a single RGBD image and produces reconstructions for each object in the scene. We treat the problem as a Bayesian inference problem over an observation described by negative samples. We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV. We use CLIP to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh. We then use SVGD to optimize for a set of samples over map weights. We can generate predicted reconstructions by taking the expected occupancy over our weights for a given location. Figure 3 shows a visual overview of our method.
In Section V-A, we explain how we leverage pre-existing mesh assets to create a prior that is robust to different poses and scales. Then, in Section V-B, we explain how we utilize the retrieval-augmented priors paradigm to retrieve relevant objects in the prior. The specific negative sampling is explained in Section V-C. Then we give the specific SVGD objective used in Section V-D.
V-A Negative Samples as Reconstruction Priors
We want to leverage existing mesh assets as our priors during Bayesian reconstruction. We define our prior as a mixture over different objects, $c_{1},...,c_{C}$. Because there is not a direct way to convert a mesh into a Hilbert map, we instead sample points $\tilde{\mathbf{x}}_{c,1},...,\tilde{\mathbf{x}}_{c,Q}\in\mathbb{R}^{3}$ around each object $c$’s mesh. We refer to these samples as the prior samples. We give them labels $\tilde{y}_{c,1},...,\tilde{y}_{c,1}\in\{-1,1\}$ determined by whether they are outside or inside the mesh. Then we simply define our prior using this data combined with a Gaussian prior over weight norm:
$P(\mathbf{w}|c):=$ $P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})$ (4)
$\propto$ $\exp(\lambda\|w\|^{2})\prod_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i})\right),$ (5)
where BCE is the same as in Equation 1.
In order to enforce pose-invariance, we first register a small stored point cloud of the object to the observed points and then transform the prior samples to this reference frame. In practice we use RANSAC [37] and the FPFH features from [38] to perform registration. In order to also have scale invariance, we do a linear scan over 10 different scales and select the the scale that resulted in the most inlier pairs from the registration.
V-B Retrieval-Augmented Priors for Hilbert Maps
Because it would be inefficient to register all meshes that are part of the prior mixture model, we propose using the retrieval-augmented prior approach introduced in Section IV. In order to determine which objects to use, we need to compute $P(c|R)$. In our case, we use CLIP [39] as a zero-shot classifier for our different objects. For each object in our prior, we store a small textual description of the object. These descriptions are then used as classes for CLIP to classify each segmented object. In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of the object and feed the cropped image for each object as the input into CLIP.
Once we have the probability of each object, we retrieve and register the stored point clouds of the top-k objects. After registration, we retrieve the prior samples corresponding to these objects to define our prior according to Equation 4.
V-C Negative Sampling
We adopt the negative sampling method introduced in [4]. The negative sampling method makes the assumption that all objects are lying on or above a planar surface. We begin by labeling the points segmented to each object as occupied for that object. Next, we perform stratified sampling along each camera ray near each object to recover a set of negatively sampled points, labeled as unoccupied. Then, we use RANSAC over points not segmented to any object to recover the flat surface all objects are resting on. This plane is used to randomly sample points in a sphere underneath each object that are near the object. We also label these points as occupied. Finally, we use grid subsampling from [40] to reduce the number of points and increase uniformity of sampled points. We refer to these points and labels as observed samples and denote them as $\{\mathbf{x}_{i}\}_{i\in[S]},\{y_{i}\}_{i\in[S]}$. The entire negative sampling process can be easily parallelized for efficient computation.
V-D SVGD Reconstruction
Once we have retrieved our prior samples and computed our observed samples, we can perform optimization-based reconstruction with SVGD. Given both sets of samples and our prior definition from Equation 4, we have the following posterior distribution:
$P(\{y_{c,i}\}|\{{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})\sum_{c\in\text{topk}}P(c|R)P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w}),$
taking the log and applying Equation 5 gives us the following objective:
$\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$
$+\sum_{c\in\text{topk}}P(c|R)\ln\left[\sum_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i}))\right)\right]$
$+\lambda\|\mathbf{w}\|^{2}+\text{const.}$.
In practice, we introduce multipliers to each term as hyperparameters during optimization, drop the constant, and use means instead of sums, creating the following objective:
$\frac{\lambda_{3}}{S}\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$ (6)
$+\frac{\lambda_{2}}{K}\sum_{c\in\text{topk}}P(c|R)\ln\left[\frac{1}{Q}\sum_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i}))\right)\right]$ (7)
$+\lambda_{\text{3}}\|\mathbf{w}\|^{2},$ (8)
where $K$ is the number of objects retrieved for the prior. This objective is used as the log of the target distribution, $\ln P(\mathbf{w})$, in Equation 2, where we also adopt the original median kernel suggested in [6]. We also opt to use SVGD in a stochastic manner, where both the observed samples and query samples are mini-batched.
From a non-probabilistic standpoint, one can interpret Equation 6 as the likelihood of the observed data, Equation 7 as the object shape prior, and Equation 8 as a regularization term.
VI Experiments
In this section, we aim to experimentally validate the following claims: (1) BRRP is more robust than deep learning methods; (2) BRRP is more accurate than methods that use uninformative priors; (3) BRRP can capture principled uncertainty. We begin by providing details on the experiments such as the baselines and metrics in Section VI-A, then we introduce results and analyses in Section VI-B
VI-A Experimental Details
BRRP Implementation: We use a set of 50 objects from the YCB dataset [41] to act as the prior for our experiments with BRRP. We implement the method in PyTorch and run the method on an NVIDIA RTX GeForce 2070 GPU.
Baselines: We compare our work against two main baselines, V-PRISM [4] and a version of PointSDF [23] that predicts occupancy and is trained on ShapeNet [42] scenes. V-PRISM is a probabilistic mapping method that uses an uninformative prior. This means that it is robust to novel objects, but doesn’t accurately reconstruct object backsides. We refer to this baseline as V-PRISM. In contrast, PointSDF is a learning based method. This means it can leverage prior information from pre-existing mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift. We refer to this baseline as PointSDF. When computing reconstruction meshes with PointSDF, a level set of $\tau=0.3$ is used.
Figure 4: Sample images of procedurally generated scenes used to evaluate BRRP. Left: a YCB scene. Right: a ShapeNet scene.
Method ShapeNet Scenes YCB Scenes Objaverse Scenes
V-PRISM [4] 0.3092 0.5003 0.4640
PointSDF [23] 0.3600 0.4601 0.3471
BRRP (ours) 0.3124 0.5277 0.4809
Table I: Intersection over union (IoU) on procedurally generated scenes from three different mesh datasets. BRRP uses a YCB prior and PointSDF is trained on ShapeNet scenes.
Figure 5: Chamfer distances (lower is better) for various methods across the procedurally generated scenes. Values are reported in centimeters. BRRP has the lowest chamfer distance on each dataset.
Figure 6: Qualitative comparison of BRRP and our baselines. PointSDF tends to predict a spherical shape for many non-spherical objects. V-PRISM can sometimes predict occupancy in portions of the scene that are not occupied. Our method is more robust and can more accurately reconstruct the scenes.
Procedurally Generated Scenes: We use the generated scenes from [4] to evaluate our method. These scenes are constructed with objects from ShapeNet [42], YCB [41], and Objaverse [43] datasets. There are 100 multi-object scenes for each mesh dataset. Each scene contains up to 10 objects. Some meshes in the Objaverse and ShapeNet scenes did not have correctly rendered textures and were instead rendered as plain white objects. Figure 4 contains two example images of these procedurally generated scenes. We also conduct an experiment on robustness where we perturb instance segmentation of the ShapeNet scenes by 2 pixels and evaluate reconstructions.
We evaluate performance on the procedurally generated scenes with two metrics: intersection over union (IoU) and chamfer distance. We refer readers to [4] for further explanation of these metrics.
Real World Scenes: In order to showcase robustness to real-world noise, we evaluate on real world scenes collected with a Kinect depth camera. In order to obtain instance segmentations, we use Grounded SAM [44] along with some depth filters. We evaluate on these real world scenes qualitatively with images of scene reconstruction and visualizing surface uncertainty.
VI-B Results
In Table I, we display the IoU results from procedurally generated scenes. The chamfer distances for the procedurally generated scenes is shown in Figure 5. The qualitative reconstructions on real world scenes can be seen in Figure 6.
Insight 1: BRRP is more accurate than a method with a uninformative prior.
As showcased in Table I, BRRP outperforms V-PRISM on each set of procedurally generated scenes. It has the highest IoU improvement when compared against V-PRISM on the YCB scenes. A similar pattern can be seen in the chamfer distances in Figure 5, where BRRP consistently outperforms V-PRISM, with the biggest gap occuring on the YCB scenes. This makes sense because BRRP uses a subset of the YCB objects as its prior.
We can see this improvement qualitatively in Figure 6. While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a portion of space that the object doesn’t occupy. BRRP is able to more accurately reconstruct these objects. The clearest example is the dark green object in the right-most scene.
Both of these quantitative and qualitative results suggest BRRP is generally more accurate than V-PRISM. It is the most accurate when evaluated on objects in its prior distribution.
Figure 7: (a) IoU of BRRP and PointSDF on ShapeNet scenes with and without shifted segmentations. Our method is more robust to segmentation shifts. (b) an example of a scene and the corresponding point cloud with shifted segmentation.
Insight 2: BRRP is more robust than a deep learning method.
While PointSDF outperformed BRRP on the ShapeNet scenes on IoU, BRRP had a lower Chamfer distance. BRRP also performed better on mesh datasets that PointSDF was not trained on. In Table I, we can see that on Objaverse scenes, where the objects were novel to both methods, BRRP performed better than PointSDF. When measuring chamfer distance, BRRP outperformed PointSDF on all datasets as shown in Figure 5. These results suggests BRRP is more robust to different object distributions than PointSDF.
Next, we evaluate robustness to slightly incorrect instance segmentations. We take the procedurally generated ShapeNet scenes and shift the segmentation over by 2 pixels. In Figure 7, we compare the IoU of BRRP and PointSDF on the scenes with and without the shift. Our method performs better on the shifted scenes compared to PointSDF.
On the real world scenes in Figure 6, BRRP is qualitatively more robust than PointSDF. PointSDF struggles with the noise associated with real world scenes as well as the novel objects. It tends to predict a spherical object on many non-spherical objects in the real world scenes. Our method on the other hand, is better able to reconstruct the objects on these scenes, including ones that are out-of-distribution for its prior.
These results suggest BRRP is more robust than PointSDF. Even though by one metric PointSDF outperforms BRRP on ShapeNet scenes, when the scenes are perturbed BRRP performs better.
Figure 8: Visualization of surface uncertainty from BRRP. Lighter areas correspond to higher uncertainty about the shape.
Insight 3: BRRP can capture principled uncertainty about object shape.
Figure 8 shows a qualitative example of uncertainty from BRRP. We measure the uncertainty by taking the variance of logits over weight particles. Our method predicts the highest uncertainty in areas of the surface that are occluded. This suggests that we can utilize surface uncertainty from BRRP in a similar way to how GPIS surface uncertainty is utilized in many grasping applications.
VII Conclusion
We introduced the concept of retrieval-augmented priors, where we retrieve relevant components of a prior distribution during Bayesian inference. We also introduced a novel Bayesian method for scene reconstruction that uses informative priors. Our method, BRRP, leveraged existing mesh datasets to build its prior and probabilistically reconstructed the scene leveraging these priors in a retrieval-augmented manner. We experimentally showed our method was more robust than a deep learning method as well as more accurate than a mapping method. Finally, showed a qualitative example of recovering principled uncertainty from our method.

### v2

@@
 Abstract
-Constructing 3D representations of object geometry is critical for many downstream manipulation tasks.
+Constructing 3D representations of object geometry is critical for many downstream robotics tasks, particularly tabletop manipulation problems.
 These representations must be built from potentially noisy partial observations.
-In this work we focus on the problem of reconstructing a multi-object scene from a single RGBD image.
-Current deep learning approaches to this problem can be brittle to noisy real world observations and out-of-distribution objects.
-Other approaches that do not rely on training data cannot accurately infer the backside of objects.
-We propose BRRP, a reconstruction method that can leverage preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction.
+In this work, we focus on the problem of reconstructing a multi-object scene from a single RGBD image, generally from a fixed camera in the scene.
+Traditional scene representation methods generally cannot infer the geometry of unobserved regions of the objects from the image.
+Attempts have been made to leverage deep learning to train on a dataset of observed objects and representations, and then generalize to new observations.
+However, this can be brittle to noisy real-world observations and objects not contained in the dataset, and cannot reason about their confidence.
+We propose BRRP, a reconstruction method that leverages preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction.
 In order to make our method more efficient, we introduce the concept of retrieval-augmented prior, where we retrieve relevant components of our prior distribution during inference.
+The prior is used to estimate the geometry of occluded portions of the in-scene objects.
 Our method produces a distribution over object shape that can be used for reconstruction or measuring uncertainty.
-We evaluate our method in both procedurally generated scenes and in real world scenes.
-We show our method is more robust than a deep learning approach while being more accurate than a method with an uninformative prior.
+We evaluate our method in both simulated scenes and in the real world.
+We demonstrate the robustness of our method against deep learning-only approaches while being more accurate than a method without an informative prior.
 I Introduction
@@
 Explicit 3D representations of the geometry of the scene are often required for the robust usage of downstream grasping and motion planning algorithms.
-These representations must be built from observations that are both noisy and only contain partial information of the scene.
+These representations must be built from observations that are both noisy and, due to occlusion, only contain partial information of the scene.
 In our case, we focus on the problem of building a 3D representation of multi-object scenes from a single RGBD camera image.
-One approach to this problem is to build some type of neural network that predicts the full geometry of an object given a partial view.
-Such approaches are able to use existing mesh datasets to more accurately infer the backside of objects.
-Unfortunately, these approaches also tend to have a number of problems when used on the real world depth cameras.
-The presence of unknown objects, significant occlusion, noisy point clouds, or inaccurate segmentation prevent these methods from being reliably deployed in less-structured environments.
-Most deep learning methods for 3D reconstruction also often lack a sense of uncertainty about the shape of the object.
-Such uncertainty has been shown to be useful in the downstream task of robotic grasping [1, 2].
-Such uncertainty can also be useful in developing safe motion plans or during active learning.
-Another common approach to building full 3D representations from a partial view is to simply build such a representation in a way that does not rely on prior information from mesh datasets.
+One approach to this problem is to train a neural network to predict the full geometry of an object given a partial view.
+Such approaches are able to use existing mesh datasets to more accurately infer the occluded backside of objects.
+Unfortunately, these approaches also tend to have a number of problems when used on real-world depth cameras.
+The presence of unknown objects, significant occlusion, noisy point clouds, or inaccurate segmentation prevents these methods from being reliably deployed in less-structured environments.
+Deep learning methods for 3D reconstruction generally lack a sense of uncertainty about the shape of the object.
+This can be particularly detrimental when an object in the scene is only partially observed and we seek to factor in the geometry of observed and occluded regions.
+Such uncertainty can enable safer and more robust operation in a range of downstream tasks, such as robot grasping [1, 2], safe motion generation and active learning.
+Another common approach to building full 3D representations from a partial view is to build a representation, solely from the observed data without considering prior information from mesh datasets.
 A common example of this is the Gaussian Process Implicit Surface [3] model.
@@
 These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution.
-However, this also means that they don’t always accurately reconstruct the backside of objects.
+However, this also means that they cannot reconstruct the unobserved backside of objects.
+Humans have the remarkable ability to infer the geometry of scenes based on prior experience, we seek to imbue robots with the same capability.
 Figure 1: Our method, BRRP, (a), (b) takes an input segmented RGBD image, and (c) retrieves objects to use as a prior, which allow it to (d) reconstruct the scene as well as (e) capture principled uncertainty about object shape.
 In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors.
-Our method is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects.
-In order to improve efficiency, we introduce the idea of a retrieval-augmented prior, inspired by [5], where we retrieve relevant components of our prior distribution based on classification results.
-We refer to our method as BRRP (pronounced ’burp’), which stands for Bayesian Reconstruction with Retrieval-augmented Priors.
-BRRP starts with an observed RGBD image with corresponding instance segmentations.
+We present Bayesian Reconstruction with Retrieval-augmented Priors (BRRP).
+BRRP is resilient to many of the pitfalls of learning-based methods while still being able to leverage an informative prior to more accurately reconstruct known objects.
+To further improve efficiency, we introduce the idea of a retrieval-augmented prior, where we retrieve relevant components of our prior distribution based on classification results.
+We begin with an observed RGBD image with corresponding instance segmentations.
 Then, we compute an identification result that predicts which objects from our database should be retrieved to use as a prior during reconstruction.
-We then negatively sample the observation and use Stein Variational Gradient Descent (SVGD) [6] to infer a posterior distribution over object shapes.
-This distribution is defined as a set of Hilbert Maps [7] for each object.
-Because we solve for a distribution over weights, we can recover principled uncertainty about the shape of each object.
-In practice we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant.
+We use this prior along with a sampled likelihood to infer a posterior distribution over object shapes.
+Because we solve for a distribution, we can recover principled uncertainty about each object’s shape.
+In practice, we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant.
 An example scene and reconstruction can be seen in Figure 1.
 We conduct experiments on BRRP in both procedurally generated scenes and in the real world.
-We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects on the generated scenes.
-We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera.
-We also qualitatively examine the uncertainty from our method’s reconstructions.
+We quantitatively show that BRRP results in accurate reconstructions and that BRRP is robust to unknown objects in the generated scenes.
+We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera, and is able to capture uncertainty within the reconstructions.
 Our three primary contributions can be summarized as follows:
@@
 II Related Works
-Explicit 3D Representations.
+3D Representations.
 There are many different ways of representing 3D geometry of a scene.
@@
 Hilbert maps [6], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space.
-Hilbert maps representations have also been extended to Bayesian Hilbert maps of various forms [9, 10, 11].
+Hilbert map representations have also been extended to Bayesian Hilbert maps of various forms [7, 8, 9].
 Neural implicit functions have also been used to represent continuous 3D geometry [10, 11, 12].
-Other representations are built using differentiable rendering and combining multiple views.
-Neural radiance fields [15] are an example of this, in which a neural network maps 3D position to density and color.
-3D Gaussian splatting [16] does a similar thing but with a set of Gaussians instead of a neural network.
-Another representation that has been studied is super quadrics [17].
+Other representations are built using differentiable rendering and combining multiple views, especially for novel-view synthesis.
+A Neural radiance field (NeRF) [13] is an example of this, in which a neural network maps 3D position to density and color.
+3D Gaussian splatting [14] tackles a similar but with a set of Gaussians instead of a neural network.
+Foundation models have also been developed for this task in [15] and applied to robotics [16].
+Other representation primitives have been studied, including super quadrics [17].
 3D Reconstruction with Deep Learning.
@@
 Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [28].
-While these work typically focus on single object scenes, other work focuses on reconstruction scenes with multiple, occluding objects.
+While these works typically focus on single object scenes, other work focuses on reconstruction scenes with multiple occluding objects.
 For example, [29] learns a reconstruction from a voxel representation with different channels to account for occlusion.
 Another method uses silhouettes to refine the initial reconstructions [30].
-In practice these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras.
+In practice, these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras.
 Inaccurate segmentation can also be a problem for many of these methods as well.
@@
 In our work, we derive our prior from pre-existing mesh datasets.
-V-PRISM [4] is another method that probabilistically maps the scene using a multiclass framing.
+V-PRISM [4] is another method that probabilistically maps the scene using a multi-class framing.
 Using Reconstructions in Manipulation.
@@
 Figure 2: (a) A hinge point feature transform induced by a set of hinge points is used by Hilbert maps [6]; (b) these maps are built by first sampling negative samples along the unoccupied portions of the camera ray.
-In order to recover the weight vector corresponding to a given depth observation, negative sampling is performed along the unoccupied portions of the depth rays.
+To recover the weight vector corresponding to a given depth observation, negative sampling is performed along the unoccupied portions of the depth rays.
 These negative samples are assigned a label of unoccupied and the points at the end of the ray are labeled as occupied.
-Then, stochastic gradient descent (SGD) is performed on the binary cross entropy of the negative samples and terminal points of the ray.
+Then, stochastic gradient descent (SGD) is performed on the binary cross entropy (BCE) of the negative samples and terminal points of the ray.
 The binary cross entropy measures the likelihood of the samples, and is defined as:
@@
 We can use this distribution to both reconstruct the scene as well as measure uncertainty.
-Stein Variational Gradient Descent (SVGD) [6] is a algorithm for variational inference that closely resembles gradient descent.
+Stein Variational Gradient Descent (SVGD) [36] is an algorithm for variational inference that closely resembles gradient descent.
 The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$.
@@
 $\hat{g}^{*}(\mathbf{x})=\frac{1}{P}\sum_{i=1}^{P}k(\mathbf{x}_{i},\mathbf{x})\nabla_{\mathbf{x_{i}}}\ln p(\mathbf{x}_{i})+\nabla_{\mathbf{x_{i}}}k(\mathbf{x},\cdot).$ (2)
-The particles can then be iteratively updated according to Equation 2:
+The particles can then be iteratively updated according to $\hat{g}^{*}$ in Equation 2 with:
 $\mathbf{x}^{(t+1)}_{i}=\mathbf{x}^{(t)}_{i}+\epsilon\hat{g}^{*}(\mathbf{x}^{(t)}_{i})$
@@
 The work has served as inspiration for an approach to affordance-prediction in [38].
-In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior distributions during Bayesian inference.
+In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior density functions during Bayesian inference.
 To motivate retrieval-augmented priors, consider the problem of Bayesian inference with a mixture model acting as the prior distribution.
@@
 Comparing to Equation 3, the expectation now replaces the true prior. Then, we can use a top-$k$ approximation for the expectation:
-$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$
+$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$ (4)
 This means that we only need to evaluate a subset of the prior distribution components.
@@
 We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV.
-We use CLIP to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh.
+We use CLIP [39] to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh.
 We then use SVGD to optimize for a set of samples over map weights.
@@
 Because it would be inefficient to register all meshes that are part of the prior mixture model, we propose using the retrieval-augmented prior approach introduced in Section IV.
-In order to determine which objects to use, we need to compute $P(c|R)$.
+In order to determine which objects to use, we need to compute $P(c|R)$ from Equation 4.
 In our case, we use CLIP [39] as a zero-shot classifier for our different objects.
@@
 These descriptions are then used as classes for CLIP to classify each segmented object.
-In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of the object and feed the cropped image for each object as the input into CLIP.
+In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of each object and feed the cropped images as input into CLIP.
 Once we have the probability of each object, we retrieve and register the stored point clouds of the top-k objects.
@@
 We refer to this baseline as V-PRISM.
-In contrast, PointSDF is a learning based method.
-This means it can leverage prior information from pre-existing mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift.
+In contrast, PointSDF is a learning-based method.
+This means it can leverage prior information from mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift.
 We refer to this baseline as PointSDF.
@@
 Figure 4: Sample images of procedurally generated scenes used to evaluate BRRP. Left: a YCB scene. Right: a ShapeNet scene.
-Method ShapeNet Scenes YCB Scenes Objaverse Scenes
-V-PRISM [4] 0.3092 0.5003 0.4640
-PointSDF [23] 0.3600 0.4601 0.3471
-BRRP (ours) 0.3124 0.5277 0.4809
+Method ShapeNet Scenes YCB Scenes Objaverse Scenes V-PRISM [4] 0.3092 0.5003 0.4640 PointSDF [23] 0.3600 0.4601 0.3471 BRRP (ours) 0.3124 0.5277 0.4809
 Table I: Intersection over union (IoU) on procedurally generated scenes from three different mesh datasets. BRRP uses a YCB prior and PointSDF is trained on ShapeNet scenes.
@@
 The qualitative reconstructions on real world scenes can be seen in Figure 6.
-Insight 1: BRRP is more accurate than a method with a uninformative prior.
+Insight 1: BRRP is more accurate than a method with an uninformative prior.
 As showcased in Table I, BRRP outperforms V-PRISM on each set of procedurally generated scenes.
-It has the highest IoU improvement when compared against V-PRISM on the YCB scenes.
-A similar pattern can be seen in the chamfer distances in Figure 5, where BRRP consistently outperforms V-PRISM, with the biggest gap occuring on the YCB scenes.
+It has the highest IoU improvement from V-PRISM on the YCB scenes.
+A similar pattern can be seen in the chamfer distances in Figure 5, where BRRP consistently outperforms V-PRISM, with the biggest gap of 0.19 (cm) occuring on the YCB scenes.
 This makes sense because BRRP uses a subset of the YCB objects as its prior.
 We can see this improvement qualitatively in Figure 6.
-While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a portion of space that the object doesn’t occupy.
+While BRRP and V-PRISM are comparable for most objects, there exist certain objects that V-PRISM predicts to occupy a large portion of space that the object doesn’t occupy.
 BRRP is able to more accurately reconstruct these objects.
-The clearest example is the dark green object in the right-most scene.
+The clearest example of this is the dark green object in the right-most scene in Figure 6.
 Both of these quantitative and qualitative results suggest BRRP is generally more accurate than V-PRISM. It is the most accurate when evaluated on objects in its prior distribution.
@@
 Insight 2: BRRP is more robust than a deep learning method.
-While PointSDF outperformed BRRP on the ShapeNet scenes on IoU, BRRP had a lower Chamfer distance.
+While PointSDF outperformed BRRP on the ShapeNet scenes on IoU (Table I), BRRP had a lower Chamfer distance (Figure 5).
 BRRP also performed better on mesh datasets that PointSDF was not trained on.
@@
 It tends to predict a spherical object on many non-spherical objects in the real world scenes.
-Our method on the other hand, is better able to reconstruct the objects on these scenes, including ones that are out-of-distribution for its prior.
+Our method on the other hand, is better able to reconstruct these scenes, including the objects that are out-of-distribution for its prior.
 These results suggest BRRP is more robust than PointSDF. Even though by one metric PointSDF outperforms BRRP on ShapeNet scenes, when the scenes are perturbed BRRP performs better.
-Figure 8: Visualization of surface uncertainty from BRRP. Lighter areas correspond to higher uncertainty about the shape.
+Figure 8: Visualization of a cylindrical Clorox container surface uncertainty from BRRP.
+Lighter areas correspond to higher uncertainty about the shape.
+We observe that the occluded back-side of the container has high uncertainty.
 Insight 3: BRRP can capture principled uncertainty about object shape.
@@
 We experimentally showed our method was more robust than a deep learning method as well as more accurate than a mapping method.
-Finally, showed a qualitative example of recovering principled uncertainty from our method.
+Finally, we showed a qualitative example of recovering principled uncertainty from our method.

### v3

@@
-Robust Bayesian Scene Reconstruction by Leveraging Retrieval-Augmented Priors
+Robust Bayesian Scene Reconstruction with Retrieval-Augmented Priors for Precise Grasping and Planning
 Abstract
-Constructing 3D representations of object geometry is critical for many downstream robotics tasks, particularly tabletop manipulation problems.
+Constructing 3D representations of object geometry is critical for many robotics tasks, particularly manipulation problems.
 These representations must be built from potentially noisy partial observations.
-In this work, we focus on the problem of reconstructing a multi-object scene from a single RGBD image, generally from a fixed camera in the scene.
-Traditional scene representation methods generally cannot infer the geometry of unobserved regions of the objects from the image.
-Attempts have been made to leverage deep learning to train on a dataset of observed objects and representations, and then generalize to new observations.
-However, this can be brittle to noisy real-world observations and objects not contained in the dataset, and cannot reason about their confidence.
+In this work, we focus on the problem of reconstructing a multi-object scene from a single RGBD image using a fixed camera.
+Traditional scene representation methods generally cannot infer the geometry of unobserved regions of the objects in the image.
+Attempts have been made to leverage deep learning to train on a dataset of known objects and representations, and then generalize to new observations.
+However, this can be brittle to noisy real-world observations and objects not contained in the dataset, and do not provide well-calibrated reconstruction confidences.
 We propose BRRP, a reconstruction method that leverages preexisting mesh datasets to build an informative prior during robust probabilistic reconstruction.
-In order to make our method more efficient, we introduce the concept of retrieval-augmented prior, where we retrieve relevant components of our prior distribution during inference.
-The prior is used to estimate the geometry of occluded portions of the in-scene objects.
-Our method produces a distribution over object shape that can be used for reconstruction or measuring uncertainty.
+We introduce the concept of a retrieval-augmented prior, where we retrieve relevant components of our prior distribution from a database of objects during inference.
+The resulting prior enables estimation of the geometry of occluded portions of the in-scene objects.
+Our method produces a distribution over object shape that can be used for reconstruction and measuring uncertainty.
 We evaluate our method in both simulated scenes and in the real world.
 We demonstrate the robustness of our method against deep learning-only approaches while being more accurate than a method without an informative prior.
+Through real-world experiments, we particularly highlight the capability of BRRP to enable successful dexterous manipulation in clutter.
 I Introduction
-The ability to construct internal representations of its operating environment is key for robot autonomy.
+To operate autonomously in novel settings, a robot must construct rich internal representations of its environment.
 These representations need to be particularly fine-grained for robotic manipulation, which often requires closely interacting with and avoiding objects.
-These interactions make it necessary for robots to develop an understanding of the geometry within their vicinity.
-Explicit 3D representations of the geometry of the scene are often required for the robust usage of downstream grasping and motion planning algorithms.
+Consider the task of robotic grasping in clutter; a robot needs to avoid unwanted collisions, while also operating in close proximity to one or more objects to make the necessary contact for the grasp.
+State of the art grasp synthesis methods often rely on motion planners to reach grasp poses in clutter [chen2024springgrasp, de2024task, matak2022planning, martin-arxiv-fpte], which need accurate 3D geometry for collision checking.
+Thus, a lack of geometric understanding can lead to unwanted collisions causing task failure or worse damage to the robot or environment.
+This is pictured in Figure 1.
+Many grasping and motion planning algorithms require explicit 3D representations of the scene’s geometry.
 These representations must be built from observations that are both noisy and, due to occlusion, only contain partial information of the scene.
-In our case, we focus on the problem of building a 3D representation of multi-object scenes from a single RGBD camera image.
+In our case, we focus on the problem of robustly building a 3D representation of multi-object scenes from a single RGBD camera image.
+Figure 1: Poor understanding of geometry can lead to unwanted collisions during robotic grasping (top), whereas robust, accurate reconstructions can enable precise collision avoidance during manipulation (bottom).
 One approach to this problem is to train a neural network to predict the full geometry of an object given a partial view.
@@
 This can be particularly detrimental when an object in the scene is only partially observed and we seek to factor in the geometry of observed and occluded regions.
-Such uncertainty can enable safer and more robust operation in a range of downstream tasks, such as robot grasping [1, 2], safe motion generation and active learning.
-Another common approach to building full 3D representations from a partial view is to build a representation, solely from the observed data without considering prior information from mesh datasets.
-A common example of this is the Gaussian Process Implicit Surface [3] model.
-A more recent example is V-PRISM [4], which probabilistically maps tabletop scenes without using prior information.
+Such uncertainty can enable safer and more robust operation in a range of downstream tasks, such as robot grasping [chen2024springgrasp, de2024task], safe motion generation and active learning.
+Another common approach to building full 3D representations from a partial view is to build a representation solely from the observed data without considering prior information from mesh datasets.
+A common example of this is the Gaussian Process Implicit Surface [dragiev2011gaussian] model.
+A more recent example is V-PRISM [wright2024vprism], which probabilistically maps tabletop scenes without using prior information.
 These reconstruction methods are more robust to unknown objects because they do not rely on any training distribution.
@@
 Humans have the remarkable ability to infer the geometry of scenes based on prior experience, we seek to imbue robots with the same capability.
-Figure 1: Our method, BRRP, (a), (b) takes an input segmented RGBD image, and (c) retrieves objects to use as a prior, which allow it to (d) reconstruct the scene as well as (e) capture principled uncertainty about object shape.
 In this work, we introduce a novel Bayesian approach for robustly reconstructing multi-object tabletop scenes by leveraging object-level shape priors.
@@
 In practice, we use a pre-existing foundation model to perform the identification and use registration to make our prior pose invariant.
-An example scene and reconstruction can be seen in Figure 1.
+Figure 2 shows an example scene reconstruction, where one of the objects (the orange juice) is decidedly out-of-distribution and our method is still able to reconstruct it according to the observed point cloud.
+Figure 2: BRRP (a), (b) takes an input segmented RGBD image, and (c) retrieves objects to use as a prior to (d) reconstruct the scene as well as (e) capture principled uncertainty about object shape, in this case, an out-of-distribution orange juice carton.
 We conduct experiments on BRRP in both procedurally generated scenes and in the real world.
@@
 We qualitatively show that BRRP is robust to noisy real world scenes collected from an RGBD camera, and is able to capture uncertainty within the reconstructions.
-Our three primary contributions can be summarized as follows:
+We also demonstrate downstream improvement with real-world grasping experiments, where the robot is able to more robustly arrive at an accurate pre-grasp pose.
+Our three primary contributions can be summarized as:
   1. 1.
@@
     A novel, robust Bayesian scene reconstruction method that utilizes prior information from existing mesh datasets
-Our paper has the following organization.
+We organize our paper as follows.
 We overview related works in Section II.
-In Section III we cover the necessary mathematical preliminaries for our method.
-In Section IV, we introduce the concept of retrieval-augmented priors.
-Our method, BRRP, is introduced in Section V.
-Then, we perform experiments in Section VI before a conclusion in Section VII
+In Section III, we cover mathematical preliminaries for our method.
+Section IV introduces retrieval-augmented priors.
+BRRP is introduced in Section V.
+Experiments are found in Section VI and a conclusion in Section VII
 II Related Works
@@
 There are many different ways of representing 3D geometry of a scene.
-In the mapping literature, techniques such as truncated signed distance functions [5] build voxelized representations of an environment.
-Hilbert maps [6], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space.
-Hilbert map representations have also been extended to Bayesian Hilbert maps of various forms [7, 8, 9].
-Neural implicit functions have also been used to represent continuous 3D geometry [10, 11, 12].
-Other representations are built using differentiable rendering and combining multiple views, especially for novel-view synthesis.
-A Neural radiance field (NeRF) [13] is an example of this, in which a neural network maps 3D position to density and color.
-3D Gaussian splatting [14] tackles a similar but with a set of Gaussians instead of a neural network.
-Foundation models have also been developed for this task in [15] and applied to robotics [16].
-Other representation primitives have been studied, including super quadrics [17].
-3D Reconstruction with Deep Learning.
+In the mapping literature, memory-intensive voxels are used as a representation of the environment.
+Hilbert maps [ramos2016hilbert], on the other hand, are a continuous occupancy map that takes the form of a linear function over some hinge point feature space.
+Hilbert map representations have also been extended to Bayesian Hilbert maps of various forms [senanayake2017bayesian, senanayake2018automorphing, zhi2019continuous].
+Neural implicit functions have also been used to represent continuous 3D geometry [park2019deepsdf, mescheder2019occupancy, sitzmann2020implicit].
+Other representations are built using differentiable rendering and combining multiple views [mildenhall2021nerf, kerbl20233d].
+Foundation models have also been developed for this task in [Wang_2024_CVPR] and applied to robotics [sim_geo_pose2024].
+We instead focus on the harder problem of reconstruction from only a single-view.
+3D Reconstruction with Deep Learning Priors.
 Many methods have been proposed as ways to leverage deep learning to reconstruct scenes or objects.
-While some methods aim to predict object shape from RGB data only [18, 19, 20, 21], we instead focus on using depth measurements during reconstruction.
-DeepSDF [10] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function.
-In the context of robotics, [22] extends DeepSDF to have uncertainty-awareness.
-Other work, such as occupancy networks [11] or PointSDF [23] try to directly predict such a latent code without inference-time optimization.
-Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [24, 25].
-Language is also used during reconstruction in [26] and [27].
-Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [28].
-While these works typically focus on single object scenes, other work focuses on reconstruction scenes with multiple occluding objects.
-For example, [29] learns a reconstruction from a voxel representation with different channels to account for occlusion.
-Another method uses silhouettes to refine the initial reconstructions [30].
+While some methods aim to predict object shape from RGB data only [xu2024instantmesh, liu2024one, engelmann2021points], we instead focus on using depth measurements during reconstruction.
+DeepSDF [park2019deepsdf] is a method to reconstruct an object by running inference-time optimization to recover a latent code for a neural implicit function.
+In the context of robotics, [liao2024uncertainty] extends DeepSDF to have uncertainty-awareness.
+Other work, such as occupancy networks [mescheder2019occupancy] or PointSDF [van2020learning] try to directly predict such a latent code without inference-time optimization.
+Deep learning has also been leveraged to learn kernels, which are used to construct a continuous signed distance function [williams2022neural, huang2023neural].
+Language is also used during reconstruction in [kasten2024point] and [cheng2023sdfusion].
+Another work uses a voxel-to-voxel variational autoencoder conditioned on bounding boxes [saund2021diverse].
+While these works typically focus on single object scenes, other work focuses on reconstructing multi-object scenes.
+For example, [agnew2021amodal] learns to reconstruct a voxel representation, leveraging different input channels to account for occlusion.
+Alternatively, work such as [iwase2024zero] aim to reconstruct the scene-wide geometry without bothering to separate each object.
+Instead, we focus on reconstructing individual objects in a manner robust to occlusion.
 In practice, these deep learning approaches can struggle to reconstruct noisy scenes with multiple, highly occluded, unknown objects on real world depth cameras.
-Inaccurate segmentation can also be a problem for many of these methods as well.
-3D Reconstruction without Deep Learning.
+These methods can also struggle when segmentation is inaccurate.
+3D Reconstruction without Deep Learning Priors.
 There are also many approaches to perform probabilistic 3D reconstruction without deep learning.
-Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [31].
-Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [3].
-While there is an extension of GPIS to a slightly more informative prior [32], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors.
-In our work, we derive our prior from pre-existing mesh datasets.
-V-PRISM [4] is another method that probabilistically maps the scene using a multi-class framing.
+Some of such methods for reconstruction use informative prior information by assuming fixed classes of objects, such as 3DP3 [gothoskar20213dp3].
+Other methods use an uninformative prior, such as Gaussian process implicit surfaces (GPIS) [dragiev2011gaussian].
+While there is an extension of GPIS to a slightly more informative prior [martens2016geometric], the only priors that can be enforced are specifically spherical, ellipsoidal, cylidrical, or planar priors.
+We instead derive our prior from pre-existing mesh data.
+V-PRISM [wright2024vprism] is another method that probabilistically maps the scene using a multi-class framing.
 Using Reconstructions in Manipulation.
 3D reconstruction methods have seen extensive use in manipulation.
-In [23], PointSDF provides collision constraints during grasping.
-PointSDF is also utilized in [33], where tactile sensors are used along with the reconstruction during grasping.
-A learning-based voxel representation is used for grasping in [34].
-Neural shape completion is also used during the anthropomorphic grasping pipeline proposed in [35].
+In [van2020learning], PointSDF provides collision constraints during grasping.
+PointSDF is also utilized in [matak2022planning], where tactile sensors are used along with the reconstruction during grasping.
+A learning-based voxel representation is used for grasping in [lundell2019robust].
+Neural shape completion is also used during the anthropomorphic grasping pipeline proposed in [hidalgo2023anthropomorphic].
 GPIS is also a common representation for manipulation applications.
-Some recent work has utilized the uncertainty from GPIS representations during grasp selection [1, 2].
+Some recent work has utilized the uncertainty from GPIS representations during grasp selection [chen2024springgrasp, de2024task].
 We believe BRRP provides principled uncertainty measurements that can similarly be utilized in downstream manipulation tasks.
@@
 III-A Hilbert Maps
-A Hilbert map [6] is a continuous occupancy map.
+A Hilbert map [ramos2016hilbert] is a continuous occupancy map.
 It represents the environment by a continuous function that is defined by a linear function of a fixed feature transform.
@@
 $m(\mathbf{x})=\sigma(\mathbf{w}^{\top}\phi(\mathbf{x})).$
-Figure 2: (a) A hinge point feature transform induced by a set of hinge points is used by Hilbert maps [6]; (b) these maps are built by first sampling negative samples along the unoccupied portions of the camera ray.
+Figure 3: (a) A hinge point feature transform induced by a set of hinge points is used by Hilbert maps [ramos2016hilbert]; (b) these maps are built by first sampling negative samples along the unoccupied portions of the camera ray.
 To recover the weight vector corresponding to a given depth observation, negative sampling is performed along the unoccupied portions of the depth rays.
@@
 $\text{BCE}(y,\mathbf{w}^{\top}\phi(\mathbf{x}))=\begin{cases}-\ln\left[\sigma\left(\mathbf{w}^{\top}\phi(\mathbf{x})\right)\right],&y=1\\ -\ln\left[1-\sigma\left(\mathbf{w}^{\top}\phi(\mathbf{x})\right)\right],&y=0\end{cases}$ (1)
-Figure 2 shows an illustration of both the hinge point feature transform and the negative samples used during Hilbert map construction.
-Hilbert maps have previously been extended to Bayesian Hilbert maps, where a distribution over weights is modeled as a multivariate Gaussian [7, 8, 9].
-There is also a multiclass variant that defines a weight matrix with Gaussian rows described in [4].
+Figure 3 shows an illustration of both the hinge point feature transform and the sampling used for creating Hilbert maps.
+Hilbert maps have previously been extended to Bayesian Hilbert maps, where a distribution over weights is modeled as a multivariate Gaussian [senanayake2017bayesian, senanayake2018automorphing, zhi2019continuous].
+There is also a multiclass variant that defines a weight matrix with Gaussian rows described in [wright2024vprism].
 In this work, we adopt the Hilbert map representation, but model the distribution over weights as a collection of particles.
@@
 III-B Stein Variational Gradient Descent
-Figure 3: Overview of BRRP method.
-We begin with a segmented RGBD image and (a) feed cropped images of each segment into CLIP to get object probabilities (Section V-B) Then, we retrieve and (b) register the the top-k objects in the prior.
-This gives us a set of registered prior samples (Section V-A).
-We also (c) compute negative samples based on the observed segmented point cloud (Section V-C).
-Finally, (d) we run SVGD optimization to recover a posterior distribution over Hilbert map weights (Section V-D).
-We can use this distribution to both reconstruct the scene as well as measure uncertainty.
-Stein Variational Gradient Descent (SVGD) [36] is an algorithm for variational inference that closely resembles gradient descent.
+Figure 4: Left: Overview of BRRP method.
+We begin with a segmented RGBD image and compute both a retrieval-augmented prior (Section V-A and Section V-B) as well as a likelihood from negative sampling (Section V-C).
+Then, we combine both in a Bayesian manner via SVGD (Section V-D).
+Right: An overview of our retrieval-augmented priors for reconstruction.
+We use CLIP to retrieve the top-$k$ relevant objects in our database (Section V-B), then we perform registration and retrieve the corresponding prior samples (Section V-A).
+Stein Variational Gradient Descent (SVGD) [liu2016stein] is an algorithm for variational inference that closely resembles gradient descent.
 The general problem of variational inference is to find a distribution $q^{*}\in\mathcal{Q}$ that is close to some target distribution $p$.
@@
 $q^{*}=\text{arg}\min_{q\in\mathcal{Q}}\mathbb{KL}(q\|p).$
-SVGD aims to iteratively transform $q$ in descent directions of the KL divergence in a $d$-dimensional reproducing kernel Hilbert space (RKHS), $\mathcal{H}^{d}$.
+SVGD aims to iteratively transform $q$ in descent directions of the KL divergence in a $d$-dimensional reproducing kernel Hilbert space, $\mathcal{H}^{d}$.
 Because this Hilbert space is a space of functions, a descent direction requires deriving the functional gradient of our KL divergence objective.
 Theorem 1
-From [36]. Let $T(\mathbf{x})=\mathbf{x}+f(\mathbf{x})$, where $f\in\mathcal{H}^{d}$ and $q_{[T]}$ is the density of random variable $\mathbf{z}=T(\mathbf{x})$ when $\mathbf{x}\sim q$. Then
+From [liu2016stein]. Let $T(\mathbf{x})=\mathbf{x}+f(\mathbf{x})$, where $f\in\mathcal{H}^{d}$ and $q_{[T]}$ is the density of random variable $\mathbf{z}=T(\mathbf{x})$ when $\mathbf{x}\sim q$. Then
 $\nabla_{f}\mathbb{KL}(q_{[T]}\|p)|_{f=0}=-g^{*}_{q,p},$
@@
 Importantly, Equation 2 only relies on the gradient of the log of $p$, which means we can perform variational inference to an unnormalized distribution.
-Such unnormalized distributions are commonplace in many Bayesian inference problems, including ours.
+Such unnormalized distributions are common in many Bayesian inference problems.
 IV Retrieval-Augmented Priors
-Retrieval-augmented generation [37] was originally introduced in the context of improving language generation.
-The work has served as inspiration for an approach to affordance-prediction in [38].
+Retrieval-augmented generation [lewis2020retrieval] was originally introduced in the context of improving language generation.
+The work has served as inspiration for an approach to affordance-prediction in [kuang2024ram].
 In our case, we draw inspiration from retrieval-augmented generation, but we use the retrieved results to improve efficiency in certain explicit formulations for prior density functions during Bayesian inference.
@@
 Given some data, we would like to infer a posterior distribution over hypotheses.
-If we have a mixture model as a prior distribution, then:
-$P(H|D)\propto P(D|H)\sum_{c=1}^{C}P(H|c).$ (3)
+Defining a mixture model prior over components $C=\{c_{1},\ldots,c_{N_{c}}\}$ gives:
+$P(H|D)\propto P(D|H)\textstyle{\sum_{c\in C}P(H|c)}.$ (3)
 If our prior distribution has a lot of components, it may be inefficient to fully evaluate.
 This could be a serious problem for algorithms like SVGD, which requires iteratively computing the gradient of both the likelihood and prior.
-Inspired by [37], the insight behind retrieval-augmented priors is to determine which subset of the prior distribution components to retrieve and use given some detection result $R$.
+Inspired by [lewis2020retrieval], the insight behind retrieval-augmented priors is to determine which subset of the prior distribution components to retrieve and use given some detection result $R$.
 Conditioning on this detection result, we have a new posterior distribution, $P(H|D,R)$.
@@
 Comparing to Equation 3, the expectation now replaces the true prior. Then, we can use a top-$k$ approximation for the expectation:
-$P(H|D,R)\propto P(D|H)\sum_{c\in\text{topk}}P(H|c)P(c|R)$ (4)
-This means that we only need to evaluate a subset of the prior distribution components.
+$P(H|D,R)\propto P(D|H)\textstyle{\sum_{c\in\text{topk}}P(H|c)P(c|R)}$ (4)
+Thus, we only evaluate a subset of the prior components.
 V The BRRP Method
@@
 We incorporate prior information on the shape of the object by leveraging retrieval-augmented priors introduced in Section IV.
-We use CLIP [39] to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh.
+We use the pretrained CLIP [radford2021learning] foundation model to determine which objects to retrieve and define our object-specific priors by a registered set of pre-computed samples from the stored mesh.
 We then use SVGD to optimize for a set of samples over map weights.
@@
 We want to leverage existing mesh assets as our priors during Bayesian reconstruction.
-We define our prior as a mixture over different objects, $c_{1},...,c_{C}$.
-Because there is not a direct way to convert a mesh into a Hilbert map, we instead sample points $\tilde{\mathbf{x}}_{c,1},...,\tilde{\mathbf{x}}_{c,Q}\in\mathbb{R}^{3}$ around each object $c$’s mesh.
+We define our prior as a mixture model with different objects serving as each component, $c_{1},...,c_{N_{c}}$.
+Because there is not a direct way to convert a mesh into a Hilbert map prior, we instead sample points $\tilde{\mathbf{x}}_{c,1},...,\tilde{\mathbf{x}}_{c,Q}\in\mathbb{R}^{3}$ around each object $c$’s mesh.
 We refer to these samples as the prior samples.
-We give them labels $\tilde{y}_{c,1},...,\tilde{y}_{c,1}\in\{-1,1\}$ determined by whether they are outside or inside the mesh.
-Then we simply define our prior using this data combined with a Gaussian prior over weight norm:
-$P(\mathbf{w}|c):=$ $P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})$ (5)
-$\propto$ $\exp(\lambda\|w\|^{2})\prod_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i})\right),$ (6)
-where BCE is the same as in Equation 1.
-In order to enforce pose-invariance, we first register a small stored point cloud of the object to the observed points and then transform the prior samples to this reference frame.
-In practice we use RANSAC [40] and the FPFH features from [41] to perform registration.
+We give them labels $\tilde{y}_{c,1},...,\tilde{y}_{c,Q}\in\{-1,1\}$ determined by whether they are outside or inside the mesh.
+Then we simply define our prior using this data and a Gaussian prior over weight norm:
+$\hat{P}(\mathbf{w}|c):=P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})$ (5)
+$\hskip 22.76219pt\propto\exp(\lambda\|w\|^{2})\prod_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i})\right).$ (6)
+In order to enforce pose-invariance for each prior component, we first register a small stored point cloud of the respective object to the observed points and then transform the prior samples to this reference frame.
+In practice we use RANSAC [fischler1981random] and the FPFH features from [rusu2009fast] to perform registration.
 In order to also have scale invariance, we do a linear scan over 10 different scales and select the the scale that resulted in the most inlier pairs from the registration.
@@
 In order to determine which objects to use, we need to compute $P(c|R)$ from Equation 4.
-In our case, we use CLIP [39] as a zero-shot classifier for our different objects.
+In our case, we use CLIP [radford2021learning] as a zero-shot classifier for our different objects.
 For each object in our prior, we store a small textual description of the object.
 These descriptions are then used as classes for CLIP to classify each segmented object.
-In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of each object and feed the cropped images as input into CLIP.
+In order to make sure CLIP knows which object we are targeting, we crop the RGB image to fit the predicted segmentation of each object.
+One could consider these cropped images to be the $R$ in our $P(c|R)$ formulation.
+We feed the cropped images as input into CLIP.
+The resulting distribution over objects predicted by CLIP becomes our $P(c|R)$ for each component.
 Once we have the probability of each object, we retrieve and register the stored point clouds of the top-$k$ objects.
-After registration, we retrieve the prior samples corresponding to these objects to define our prior according to Equation 5.
+After registration, we retrieve the prior samples corresponding to these objects to define our top-$k$ component prior approximation according to Equation 5.
 V-C Negative Sampling
-We adopt the negative sampling method introduced in [4].
+We adopt the negative sampling method introduced in [wright2024vprism].
 The negative sampling method makes the assumption that all objects are lying on or above a planar surface.
@@
 We also label these points as occupied.
-Finally, we use grid subsampling from [42] to reduce the number of points and increase uniformity of sampled points.
+Finally, we use grid subsampling from [thomas2019learning] to reduce the number of points and increase uniformity of sampled points.
 We refer to these points and labels as observed samples and denote them as $\{\mathbf{x}_{i}\}_{i\in[S]},\{y_{i}\}_{i\in[S]}$.
 The entire negative sampling process can be easily parallelized for efficient computation.
+Figure 5: Sample images of procedurally generated scenes used to evaluate BRRP. Left: a YCB scene. Right: a ShapeNet scene.
 V-D SVGD Reconstruction
 Once we have retrieved our prior samples and computed our observed samples, we can perform optimization-based reconstruction with SVGD.
-Given both sets of samples and our prior definition from Equation 5, we have the following posterior distribution:
-$P(\{y_{c,i}\}|\{{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})\sum_{c\in\text{topk}}P(c|R)P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w}),$
-taking the log and applying Equation 6 gives us the following objective:
-$\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$
-$+\sum_{c\in\text{topk}}P(c|R)\ln\left[\sum_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i}))\right)\right]$
-$+\lambda\|\mathbf{w}\|^{2}+\text{const.}$.
-In practice, we introduce multipliers to each term as hyperparameters during optimization, drop the constant, and use means instead of sums, creating the following objective:
-$\frac{\lambda_{3}}{S}\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$ (7)
+Given both sets of samples and our prior definition from Equation 5, we have the following posterior distribution,
+$P(\{y_{c,i}\}|\{{\mathbf{x}}_{c,i}\},\mathbf{w})P(\mathbf{w})\sum_{c\in\text{topk}}P(c|R)P(\{\tilde{y}_{c,i}\}|\{\tilde{\mathbf{x}}_{c,i}\},\mathbf{w}).$
+Taking the log and applying Equation 6 gives the objective:
+$\ell=\frac{\lambda_{\text{{1}}}}{S}\sum_{i=1}^{S}\text{BCE}(y_{i},\mathbf{w}^{\top}\phi(\mathbf{x}_{i}))$ (7)
 $+\frac{\lambda_{2}}{K}\sum_{c\in\text{topk}}P(c|R)\ln\left[\frac{1}{Q}\sum_{i=1}^{Q}\exp\left(\text{BCE}(\tilde{y}_{c,i},\mathbf{w}^{\top}\phi(\tilde{\mathbf{x}}_{c,i}))\right)\right]$ (8)
@@
 where $K$ is the number of objects retrieved for the prior.
-This objective is used as the log of the target distribution, $\ln P(\mathbf{w})$, in Equation 2, where we also adopt the original median kernel suggested in [36].
+For practical reasons, we introduce multipliers to each term as hyperparameters during optimization and use means instead of sums for the first two terms.
+This objective is used as the log of the target distribution, $\ln P(\mathbf{w})$, in Equation 2, where we also adopt the original median kernel suggested in [liu2016stein].
 We also opt to use SVGD in a stochastic manner, where both the observed samples and query samples are mini-batched.
-From a non-probabilistic standpoint, one can interpret Equation 7 as the likelihood of the observed data, Equation 8 as the object shape prior, and Equation 9 as a regularization term.
+From a non-probabilistic standpoint, one can interpret Equation 7 as the likelihood of the observed data, Equation 8 as the object shape prior, and Equation 9 as regularization.
 VI Experiments
 In this section, we aim to experimentally validate the following claims: (1) BRRP is more robust than deep learning methods; (2) BRRP is more accurate than methods that use uninformative priors; (3) BRRP can capture principled uncertainty.
-We begin by providing details on the experiments such as the baselines and metrics in Section VI-A, then we introduce results and analyses in Section VI-B
+(4) BRRP’s robustness improves the downstream success of grasping.
+First, we provide details on the experiments, such as the baselines and metrics in Section VI-A, then we show then show the results and analyses in Section VI-B.
+Refer to caption Refer to caption
+Figure 6: Results pertaining to the speed of BRRP on an NVIDIA RTX 4070 GPU.
+Left: Ablation on $k$ for YCB scenes highlights the accuracy-speed tradeoff in retrieval-augmented priors.
+We use $k=3$ for all other experiments.
+Right: Timing breakdown of BRRP on a real-world scene with 4 objects.
 VI-A Experimental Details
-BRRP Implementation: We use a set of 50 objects from the YCB dataset [43] to act as the prior for our experiments with BRRP.
-We implement the method in PyTorch and run the method on an NVIDIA RTX GeForce 2070 GPU.
-Baselines: We compare our work against two main baselines, V-PRISM [4] and a version of PointSDF [23] that predicts occupancy and is trained on ShapeNet [44] scenes.
+BRRP Implementation: We use a set of 50 objects from the YCB dataset [calli2015ycb] to act as the prior for our experiments with BRRP.
+We implement the method in PyTorch.
+We also display results for the speed of running BRRP on an NVIDIA RTX 4070 GPU in Figure 6.
+We use $k=3$ for all experiments, but note that our method’s performance is fairly robust to the $k$ parameter (Figure 6).
+Baselines: We compare our work against two main baselines, V-PRISM [wright2024vprism] and a version of PointSDF [van2020learning] that predicts occupancy and is trained on ShapeNet [chang2015shapenet] scenes.
 V-PRISM is a probabilistic mapping method that uses an uninformative prior.
@@
 This means it can leverage prior information from mesh datasets to accurately reconstruct the backside, but can suffer in performance under significant distributional shift.
-We refer to this baseline as PointSDF.
-When computing reconstruction meshes with PointSDF, a level set of $\tau=0.3$ is used.
-Figure 4: Sample images of procedurally generated scenes used to evaluate BRRP. Left: a YCB scene. Right: a ShapeNet scene.
-Method ShapeNet Scenes YCB Scenes Objaverse Scenes V-PRISM [4] 0.3092 0.5003 0.4640 PointSDF [23] 0.3600 0.4601 0.3471 BRRP (ours) 0.3124 0.5277 0.4809
-Table I: Intersection over union (IoU) on procedurally generated scenes from three different mesh datasets. BRRP uses a YCB prior and PointSDF is trained on ShapeNet scenes.
+We train two versions of PointSDF.
+One is trained on scenes containing a wide variety of Shapenet objects, and another trained only on scenes containing the 50 YCB objects in our BRRP prior.
+We refer to this baseline as PointSDF, and generally use the ShapeNet version when the training dataset is not mentioned.
+When reconstructing meshes with PointSDF, a level set of $\tau=0.3$ is used.
+Method ShapeNet Scenes (IoU $\uparrow$) YCB Scenes (IoU $\uparrow$) V-PRISM [wright2024vprism] 0.3092 0.5003 PointSDF (ShapeNet) [van2020learning] 0.3600 0.4601 PointSDF (YCB)[van2020learning] 0.2426 0.5088 BRRP (ours) 0.3124 0.5277
+Table I: Intersection over union (IoU) on procedurally generated scenes from two different mesh datasets.
+BRRP uses a YCB prior and there are two different PointSDF versions, trained on ShapeNet and YCB scenes respectively.
+BRRP outperforms V-PRISM, which has an uniformative prior, whereas BRRP is more robust to distributional shift than PointSDF.
 Figure 7: Chamfer distances (lower is better) for various methods across the procedurally generated scenes. Values are reported in centimeters. BRRP has the lowest chamfer distance on each dataset.
@@
 Our method is more robust and can more accurately reconstruct the scenes.
-Procedurally Generated Scenes: We use the generated scenes from [4] to evaluate our method.
-These scenes are constructed with objects from ShapeNet [44], YCB [43], and Objaverse [45] datasets.
+Procedurally Generated Scenes: We use the generated scenes from [wright2024vprism] to evaluate our method.
+These scenes are constructed with objects from ShapeNet [chang2015shapenet], YCB [calli2015ycb], and Objaverse [deitke2023objaverse] datasets.
 There are 100 multi-object scenes for each mesh dataset.
@@
 Figure 5 contains two example images of these procedurally generated scenes.
-We also conduct an experiment on robustness where we perturb instance segmentation of the ShapeNet scenes by 2 pixels and evaluate reconstructions.
-We evaluate performance on the procedurally generated scenes with two metrics: intersection over union (IoU) and chamfer distance. We refer readers to [4] for further explanation of these metrics.
+We use the ground-truth masks for experiments on procedural scenes.
+We also conduct an experiment on robustness where we perturb the masks of the ShapeNet scenes by two pixels and evaluate reconstructions.
+We evaluate performance on the procedurally generated scenes with two metrics: intersection over union (IoU) and chamfer distance.
+We refer readers to [wright2024vprism] for further explanation of these metrics.
 Real World Scenes: In order to showcase robustness to real-world noise, we evaluate on real world scenes collected with a Kinect depth camera.
-In order to obtain instance segmentations, we use Grounded SAM [46] along with some depth filters.
+In order to obtain instance segmentations, we use Grounded SAM [ren2024grounded] along with some depth filters.
+We include timing of performing this segmentation in the real-world timing breakdown in Figure 6.
 We evaluate on these real world scenes qualitatively with images of scene reconstruction and visualizing surface uncertainty.
-VI-B Results
+Grasping Experiment: To understand the downstream impact of our method, we assess its effectiveness on the problem of dexterous grasping from a fixed RGB-D camera.
+We place an object on the table, segment it, reconstruct its mesh, plan a grasp trajectory, and execute it open-loop.
+Figure 9 shows the 10 objects used.
+Figure 9: Objects and the robot used to assess the impact of our method on dexterous grasping.
+For segmentation, we use Grounded SAM [ren2024grounded] with the prompt object that can be picked up with one hand.
+Meshes are reconstructed with marching cubes on either BRRP or PointSDF reconstructions.
+We implement a grasp planner [martin-arxiv-fpte] which takes a segmented point cloud as input and moves the robot to a pregrasp configuration that is collision-free according to the reconstruction while also avoiding obstacles according to the scene reconstruction.
+The robot then closes the hand and lifts the arm.
+An attempt is successful if the object is lifted.
+We use KUKA LBR4 with Allegro hand in the experiment.
+We place each of the 10 objects across 5 different poses on the table, resulting in 50 executed grasp trajectories for each method.
+We select stable poses where objects are reachable by the robot, visible to the camera, and remain stationary, while ensuring diversity in positions and orientations.
+We report the grasp success rate per method across all 50 trajectories.
+VI-B Results & Discussion
 In Table I, we display the IoU results from procedurally generated scenes.
@@
 Insight 1: BRRP is more accurate than a method with an uninformative prior.
-As showcased in Table I, BRRP outperforms V-PRISM on each set of procedurally generated scenes.
+As showcased in Table I, BRRP outperforms V-PRISM on both sets of procedurally generated scenes.
 It has the highest IoU improvement from V-PRISM on the YCB scenes.
@@
 The clearest example of this is the dark green object in the right-most scene in Figure 8.
-Both of these quantitative and qualitative results suggest BRRP is generally more accurate than V-PRISM. It is the most accurate when evaluated on objects in its prior distribution.
-Figure 7: (a) IoU of BRRP and PointSDF on ShapeNet scenes with and without shifted segmentations.
+Both of these quantitative and qualitative results suggest BRRP is generally more accurate than V-PRISM.
+It is the most accurate when evaluated on objects in its prior distribution.
+Figure 10: (left) IoU of BRRP and PointSDF on ShapeNet scenes with and without shifted segmentations.
 Our method is more robust to segmentation shifts.
-(b) An example of a scene and the corresponding point cloud with shifted segmentation.
+(right) An example of a scene and the corresponding point cloud with shifted segmentation.
 Insight 2: BRRP is more robust than a deep learning method.
-While PointSDF outperformed BRRP on the ShapeNet scenes on IoU (Table I), BRRP had a lower Chamfer distance (Figure 5).
-BRRP also performed better on mesh datasets that PointSDF was not trained on.
-In Table I, we can see that on Objaverse scenes, where the objects were novel to both methods, BRRP performed better than PointSDF.
-When measuring chamfer distance, BRRP outperformed PointSDF on all datasets as shown in Figure 5.
+While the ShapeNet version of PointSDF outperformed BRRP on the ShapeNet scenes on IoU (Table I), BRRP had a lower Chamfer distance (Figure 7).
+BRRP also outperformed the YCB version of PointSDF on both YCB and ShapeNet (Table I).
+Generally, BRRP performed significantly better than each PointSDF version on mesh datasets the specific version was not trained on.
+When measuring chamfer distance, BRRP outperformed the ShapeNet version of PointSDF on all datasets as shown in Figure 7, including the Objaverse dataset, where objects were out-of-distribution for both methods.
 These results suggests BRRP is more robust to different object distributions than PointSDF.
-Next, we evaluate robustness to slightly incorrect instance segmentations.
+Next, we evaluate robustness to slightly incorrect image segmentations.
 We take the procedurally generated ShapeNet scenes and shift the segmentation over by 2 pixels.
-In Figure 7, we compare the IoU of BRRP and PointSDF on the scenes with and without the shift.
+In Figure 10, we compare the IoU of BRRP and ShapeNet version of PointSDF on the scenes with and without the shift.
 Our method performs better on the shifted scenes compared to PointSDF.
-On the real world scenes in Figure 6, BRRP is qualitatively more robust than PointSDF.
-PointSDF struggles with the noise associated with real world scenes as well as the novel objects.
+On the real-world scenes in Figure 8, BRRP is qualitatively more robust than PointSDF.
+PointSDF struggles with the noise inherent in real world scenes as well as novel objects.
 It tends to predict a spherical object on many non-spherical objects in the real world scenes.
-Our method on the other hand, is better able to reconstruct these scenes, including the objects that are out-of-distribution for its prior.
-These results suggest BRRP is more robust than PointSDF. Even though by one metric PointSDF outperforms BRRP on ShapeNet scenes, when the scenes are perturbed BRRP performs better.
+Our method is better able to reconstruct these scenes, including the objects that are out-of-distribution for its prior.
+These results suggest BRRP is more robust than PointSDF.
+Even though by one metric the ShapeNet-trained PointSDF outperforms BRRP on ShapeNet scenes, when the scenes are perturbed, BRRP performs better.
 Figure 11: Visualization of a cylindrical Clorox container surface uncertainty from BRRP.
@@
 We observe that the occluded back-side of the container has high uncertainty.
+Figure 12: BRRP builds precise representations, which are critical for downstream dexterous manipulation.
+Here, the water (out-of-distribution) is particularly large, and the dexterous hand needs to arrive at a pre-grasp pose that is very close to the surface of the bottle, all the while requiring the fingers to avoid colliding.
 Insight 3: BRRP can capture principled uncertainty about object shape.
@@
 This suggests that we can utilize surface uncertainty from BRRP in a similar way to how GPIS surface uncertainty is utilized in many grasping applications.
+Insight 4: BRRP’s robust reconstructions improve the downstream success of dexterous grasping.
+The success rate of BRRP and the PointSDF baseline is 64% and 40% respectively on the downstream grasping experiment.
+The difference in success rates stems from the modeling of the obstacles, consisting of the reconstructed object mesh and the table (fixed).
+The collision modeling directly impacts how close the robot can get to the object before executing the grasp.
+This directly affects whether the object will get knocked over or bumped into as the gripper moves to the pre-grasp configuration.
+BRRP exhibits higher accuracy and robustness in reconstruction, which creates an over 50% increase in grasp success rates.
+Figure 12 illustrates an example and highlights the impact of BRRP in significantly improving downstream grasp success.
+While the difference in BRRP and PointSDF grasp success can be easily attributed to reconstruction quality, we hypothesize the 36% of grasp failures with BRRP stem from a mixture of reconstruction, grasp selection, and motion planning inaccuracies.
+Limitations: As shown in Figure 6, the registration portion of our method grows with the number of objects, which can be resolved by fast learning-based registration methods.
+While our method is robust to viewpoint, occlusion, and scale; it is still limited by the quality of segmentation, depth accuracy, and its prior database.
+Advances in segmentation methods will further improve BRRP performance.
+While BRRP can robustly handle OOD data, the reconstruction and uncertainty remain marginally biased towards the in-distribution dataset, which is common for data-driven approaches.
+This can be seen in Figure 2 and Figure 8, and can be mitigated with a larger, more diverse prior database.
 VII Conclusion
@@
 Our method, BRRP, leveraged existing mesh datasets to build its prior and probabilistically reconstructed the scene leveraging these priors in a retrieval-augmented manner.
-We experimentally showed our method was more robust than a deep learning method as well as more accurate than a mapping method.
-Finally, we showed a qualitative example of recovering principled uncertainty from our method.
+We showed our method was more robust than a deep learning method as well as more accurate than a mapping method.
+We also showed a qualitative example of recovering principled uncertainty from our method.
+Finally, we demonstrated that our robust reconstructions improve downstream performance, where BRRP reconstructions enable precise collision avoidance as the robot moves to execute grasps.